# Partial Condition Numbers for Double Saddle Point Problems

### v1

@@ 1 Introduction
$A\in{\mathbb{R}}^{n\times n},$ $B\in{\mathbb{R}}^{m\times n},$ $C\in{\mathbb{R}}^{p\times m},$ $D\in{\mathbb{R}}^{m\times m},$ $E\in{\mathbb{R}}^{p\times p},$ $\mathbf{b}\in{\mathbb{R}}^{\bm{l}\times\bm{l}}$ and $\bm{l}=n+m+p.$ Then the linear system of the form (1.1) is known as the double saddle point problem (DSPP). We refer $\mathfrak{B}$ as the double saddle point matrix. Conditions on the invetability for the matrix $\mathfrak{B}$ have been studied in [1, 2]. To ensure a unique solution to (1.1), throughout the paper, we assume that $\mathfrak{B}$ is nonsingular.
@@
Linear systems of the form (1.1) emerge in numerous computational science and technological applications, such as in finite element discretizations of coupled Stokes-Darcy flow equations [3], the equality constrained indefinite least square problem (EILS) [4], computational fluid dynamics [5], liquid crystal director modeling [6], PDE-constraint optimization problems [7], and so on. Due to the versatile applicability of the DSPP (1.1), their numerical solutions have garnered significant interest and have been extensively investigated in recent years. Several iterative methods and preconditioning techniques have been developed in recent times for solving DSPPs; see [2, 8, 9, 10] and reference therein.
@@
Perturbation analysis is one of the most useful tool in numerical analysis for assessing the sensitivity and stability of the solution of a problem obtained through numerical methods [11]. The condition number (CN) quantifies the worst-case sensitivity of a solution to a small perturbation in the input data matrices. Additionally, CNs are vital in providing first-order error estimates for the relative error of a numerical solution. The general framework of CNs was first introduced by Rice [12]. It primarily focuses on the normwise CN (NCN), where norms are used to quantify both the input perturbations and the resulting errors in the output data. However, a significant limitation of NCN is its inability to take into account the inherent structure of poorly scaled or sparse input data. To overcome this limitation, the mixed CN (MCN) and componentwise CN (CCN) have gained increasing attention in the literature [13, 14, 15]. The former quantifies input perturbations componentwise and output errors using norms, while the later measures both input perturbations and output errors componentwise.
Perturbation analysis and CNs for standard saddle point problems (SPPs) have been extensively studied in the literature; see [16, 17, 18, 19]. Although, these studies fail to leverage the three-by-three block structure of the coefficient matrix $\mathfrak{B}$. Moreover, in many applications, the solution components $\bm{x}$, $\bm{y}$ and $\bm{z}$ of $\bm{w}$ represent different physical entities. For instance, in PDE-constrained optimization problems [7], $\bm{x}$ represents the desired state, $\bm{z}$ corresponds to the control variable, and $\bm{y}$ denotes the Lagrange multipliers. However, the sensitivity of each solution component can vary; hence, analyzing the sensitivity of the individual solution components $\bm{x},$ $\bm{y}$ and $\bm{z}$ or each solution component of $\bm{w}$ becomes crucial [20]. The traditional CNs lack the ability to reveal the conditioning of a specific part of the solution. To tackle this situation, CN of a linear function of the solution has been investigated in the literature, by introducing a matrix $\mathrm{L}$. This matrix enables us to select any particular solution component. The CN of a linear function of the solution is referred to as the partial CN (or the projected CN). For the first time, this class of CN was investigated in [20] for the system of linear equations and later extensively studied for various problems in recent years, for instance, in linear least squares problems [21], weighted least squares problems [22], the indefinite least squares problems [23, 24], total least squares problems [25], generalized SPPs [19].
@@
In recent years, the structured CNs of various problems have been studied, emphasizing the preservation of the linear structure of the original matrices in the perturbation matrices; see, for example [26, 27, 28, 19]. The block matrices $A$, $D$ and $E$ often exhibit particular linear structures in various applications; see [7, 6]. This makes it compelling to explore the structured partial CNs for the DSPP (1.1) by preserving the linear structures of the diagonal block matrices to their corresponding perturbation matrices.
@@
  • In this work, we explore a general form of partial CN, refer as unified partial CN, which has a versatile nature and provides a comprehensive framework encompassing the partial NCN, MCN and CCN of the solution of the DSPP.
@@ 2.2 Preliminaries
$\left\|\xi^{\ddagger}\odot(\Upsilon(\bm{v}+{\Delta}\bm{v})-\Upsilon(\bm{v}))\right\|_{\gamma}\leq\mathfrak{K}_{\Upsilon}(\bm{v})\|\chi^{\ddagger}\odot{\Delta}\bm{v}\|_{\tau}+\mathcal{O}(\|\chi^{\ddagger}\odot{\Delta}\bm{v}\|^{2}_{\tau}).$ (2.11)
@@ 3 Partial Unified CNs for the DSPP
In the following theorem, we provide compact and closed-form expression for the partial unified CN.
@@
Notably, the equivalent expression of $\widehat{\mathfrak{K}}^{(2)}_{\Phi}(\mathbf{H},{\bf b};\mathrm{L})$ in (3.7) is free of computationally expensive Kronecker products. Moreover, the matrices in (3.6) and (3.7) have dimensions $k\times(\bm{l+s})$ and $k\times k$, respectively. Hence, the expression in (3.7) significantly reduces the storage requirements.
Remark 4.
@@
In the next result, we provide an easily computable upper bound for the partial CN $\mathfrak{K}^{(2)}_{\Phi}(\mathbf{H},{\bf b};\mathrm{L}).$
@@
In the following result, we provide sharp upper bounds for the partial MCN and CCN obtained in Theorem 3.
@@ 4 Structured partial CNs
Consider three subspaces $\mathbb{S}_{1}\subseteq{\mathbb{R}}^{n\times n}$, $\mathbb{S}_{2}\subseteq{\mathbb{R}}^{m\times m}$ and $\mathbb{S}_{3}\subseteq{\mathbb{R}}^{p\times p}$ consisting of three distinct linear structured matrices, such as symmetric. Suppose that the corresponding dimensions of the linear subspaces are $s,$ $r$ and $q,$ respectively. Let $A\in\mathbb{S}_{1},$ $D\in\mathbb{S}_{2}$ and $E\in\mathbb{S}_{3},$ then according to [33, 34, 26], there exist unique generating vectors $\bm{a}\in{\mathbb{R}}^{s}$ $\bm{d}\in{\mathbb{R}}^{r}$ and $\bm{e}\in{\mathbb{R}}^{q}$ such that
@@ 6 Numerical experiments
In this part, we present some numerical examples to verify the reliability of the derived the partial NCN, MCN and CCN and their upper bounds for the DSPP. Additionally, we demonstrate their effectiveness in providing tight upper bounds for the relative forward error in solving the DSPP.
@@
respectively. Define the following quantities:
@@
Thus from (2.11), $\bm{\epsilon}_{1}\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}),$ $\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ and $\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ can be employed to estimate the relative forward errors $\bm{r}_{k},$ $\bm{r}_{m}$ and $\bm{r}_{c},$ respectively. We select the matrix $\mathrm{L}$ as
@@
where $J=\frac{1}{(q+1)^{2}}\,\mathtt{tridiag}(-1,2,-1)\in{\mathbb{R}}^{q\times q},$ $Z=\frac{1}{q+1}\,\mathtt{tridiag}(0,1,-1)\in{\mathbb{R}}^{q\times q}$ and $Y=\mathrm{diag}(1,q+1,\ldots,q^{2}-q+1)\in{\mathbb{R}}^{q\times q}.$ Further, we take $D=I_{m}$ and $E=I_{p}.$ The notation $\mathtt{tridiag}(d_{1},d_{2},d_{3})\in{\mathbb{R}}^{q\times q}$ denotes the tridiagonal matrix with subdiagonal entries $d_{1}$, diagonal entries $d_{2}$ and superdiagonal entries $d_{3}$. Here, the dimension of the coefficient matrix $\mathfrak{B}$ of the DSPP is ${\bm{l}}=4q^{2}.$ We take ${\bf b}=\tt{randn}(n,1)\in{\mathbb{R}}^{\bm{l}}$. Further, we consider $\Psi=\|\mathfrak{B}\|_{F}$ and ${\bm{\chi}}=\|{\bf b}\|_{2}.$ We use Theorems 2, 3 to compute partial CNs and Corollaries 3.1 and 3.2 for their upper bounds. The numerical results for different choices for $\mathrm{L}$ and $q=4:2:16$ are presented in Tables 1-4.
Table 1: Comparison of the NCN, MCN, and CCN, and their upper bounds, with the corresponding relative errors for $\mathrm{L}=\mathrm{L}_{0}$ for Example 6.1.
$q$ $\bm{r}_{k}$ $\mathfrak{K}_{2}$ $\mathfrak{K}_{2}^{U}$ $\bm{r}_{m}$ $\mathfrak{K}_{m}$ $\mathfrak{K}_{m}^{U}$ $\bm{r}_{c}$ $\mathfrak{K}_{c}$ $\mathfrak{K}_{c}^{U}$
@@
Table 2: Comparison of the NCN, MCN, and CCN, and their upper bounds, with the corresponding relative errors for $\mathrm{L}=\mathrm{L}_{n}$ for Example 6.1.
$q$ $\bm{r}_{k}$ $\mathfrak{K}_{2}$ $\mathfrak{K}_{2}^{U}$ $\bm{r}_{m}$ $\mathfrak{K}_{m}$ $\mathfrak{K}_{m}^{U}$ $\bm{r}_{c}$ $\mathfrak{K}_{c}$ $\mathfrak{K}_{c}^{U}$
@@
Table 3: Comparison of the NCN, MCN, and CCN, and their upper bounds, with the corresponding relative errors for $\mathrm{L}=\mathrm{L}_{m}$ for Example 6.1.
$q$ $\bm{r}_{k}$ $\mathfrak{K}_{2}$ $\mathfrak{K}_{2}^{U}$ $\bm{r}_{m}$ $\mathfrak{K}_{m}$ $\mathfrak{K}_{m}^{U}$ $\bm{r}_{c}$ $\mathfrak{K}_{c}$ $\mathfrak{K}_{c}^{U}$
@@
Table 4: Comparison of the NCN, MCN, and CCN, and their upper bounds, with the corresponding relative errors for $\mathrm{L}=\mathrm{L}_{p}$ for Example 6.1.
$q$ $\bm{r}_{k}$ $\mathfrak{K}_{2}$ $\mathfrak{K}_{2}^{U}$ $\bm{r}_{m}$ $\mathfrak{K}_{m}$ $\mathfrak{K}_{m}^{U}$ $\bm{r}_{c}$ $\mathfrak{K}_{c}$ $\mathfrak{K}_{c}^{U}$
@@
$\mathfrak{K}_{2}=\bm{\epsilon}_{1}\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),\leavevmode\nobreak\ \mathfrak{K}_{2}^{U}=\bm{\epsilon}_{1}\mathfrak{K}^{(2),u}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),\leavevmode\nobreak\ \mathfrak{K}_{m}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),$
@@
$\mathfrak{K}_{m}^{U}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty,u}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),\leavevmode\nobreak\ \mathfrak{K}_{c}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}),\leavevmode\nobreak\ \mathfrak{K}_{c}^{U}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty,u}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}_{m}).$

### v2

@@ 1 Introduction
$A\in{\mathbb{R}}^{n\times n},$ $B\in{\mathbb{R}}^{m\times n},$ $C\in{\mathbb{R}}^{p\times m},$ $D\in{\mathbb{R}}^{m\times m},$ $E\in{\mathbb{R}}^{p\times p},$ $\mathbf{b}\in{\mathbb{R}}^{\bm{l}}$ and $\bm{l}=n+m+p.$ Then the linear system of the form (1.1) is known as the double saddle point problem (DSPP). We refer $\mathfrak{B}$ as the double saddle point matrix. Conditions on the invetability for the matrix $\mathfrak{B}$ have been studied in [1, 2]. To ensure a unique solution to (1.1), throughout the paper, we assume that $\mathfrak{B}$ is nonsingular.
@@
Linear systems of the form (1.1) emerge in numerous computational science and technological applications, such as in finite element discretizations of coupled Stokes-Darcy flow equations [3], the equality constrained indefinite least square problem (EILS) [4], computational fluid dynamics [5], liquid crystal director modeling [6], PDE-constraint optimization problems [7], and so on. Due to the versatile applicability of the DSPP (1.1), their numerical solutions have garnered significant interest and have been extensively investigated in recent years. Several iterative methods and preconditioning techniques have been developed in recent times for solving DSPPs; see [2, 8, 9, 10] and references therein.
@@
The coefficient matrix of a DSPP is typically obtained from measurements and is therefore subject to some measurement errors, as well as subject to the rounding and truncation errors introduced due to the finite precision of the computers. As a result of these inevitable perturbations, we work with a matrix that is slightly different from the original one. Since such errors are often unavoidable, it is important to investigate how small perturbations in the input matrices influence the resulting solution. Perturbation analysis is one of the most useful tool in numerical analysis for assessing the sensitivity and stability of the solution of a problem obtained through numerical methods [11]. The condition number (CN) quantifies the worst-case sensitivity of a solution to a small perturbation in the input data matrices. Additionally, CNs are vital in providing first-order error estimates for the relative error of a numerical solution. The general framework of CNs was first introduced by Rice [12]. It primarily focuses on the normwise CN (NCN), where norms are used to quantify both the input perturbations and the resulting errors in the output data. However, a significant limitation of NCN is its inability to take into account the inherent structure of poorly scaled or sparse input data. To overcome this limitation, the mixed CN (MCN) and componentwise CN (CCN) have gained increasing attention in the literature [13, 14, 15]. The former quantifies input perturbations componentwise and output errors using norms, while the later measures both input perturbations and output errors componentwise.
Perturbation analysis, CNs, and structured backward error analysis for standard saddle point problems (SPPs) of the following form:
$\left[\begin{matrix}F&G^{{\bm{\top}}}\\ H&K\end{matrix}\right]\left[\begin{matrix}u\\ p\end{matrix}\right]=\left[\begin{matrix}f\\ g\end{matrix}\right]$ (1.3)
have been extensively studied in the literature. For example, the NCN for the solution of the SPP (1.3), under the assumptions $F=F^{{\bm{\top}}}$, $G=H$, and $K=\bf 0$, was initially investigated in [16]. In [17], explicit expressions for the NCN, MCN, and CCN were derived for the solution vector $[u^{{\bm{\top}}},\,p^{{\bm{\top}}}]^{{\bm{\top}}}$, as well as for its components $u$ and $p$, under the same structural assumptions. Building on this, Xu and Li [18] further examined the NCN and established perturbation bounds for the combined solution vector $[u^{\bm{\top}},\,p^{\bm{\top}}]^{\bm{\top}}$. Subsequently, Meng and Li [19] extended this line of research by analyzing both the MCN and CCN for the full solution and its individual components $u$ and $p$. More recently, structured CNs and backward error analysis for the SPP (1.3) have been explored in [20, 21], particularly when the coefficient matrices exhibit linear structures such as Toeplitz or symmetry. However, these studies fail to leverage the special three-by-three block structure of the coefficient matrix $\mathfrak{B}$, and the CNs for the DSPP (1.1) remain unexplored and still require further detailed investigation.
@@
In recent years, the structured CNs of various problems have been studied, emphasizing the preservation of the linear structure of the original matrices in the perturbation matrices; see, for example [22, 23, 24, 20]. The block matrices $A$, $D$ and $E$ often exhibit particular linear structures in various applications; see [7, 6]. This makes it compelling to explore the structured CNs for the DSPP (1.1) by preserving the linear structures of the diagonal block matrices to their corresponding perturbation matrices.
In many applications, the components $\bm{x}$, $\bm{y}$, and $\bm{z}$ of the solution vector $\bm{w}$ correspond to distinct physical quantities. For example, in PDE-constrained optimization problems [7], $\bm{x}$ typically represents the desired state, $\bm{z}$ the control variable, and $\bm{y}$ the Lagrange multipliers. Since each component may respond differently to perturbations, it is essential to analyze the sensitivity of individual components rather than the full solution vector [25]. Traditional CNs, which capture the overall sensitivity, lack the ability to reveal the conditioning of a specific part of the solution. To overcome this limitation, the concept of a partial or projected CN was introduced, which quantifies the sensitivity of a linear function of the solution using a selection matrix $\mathrm{L}$. This matrix allows one to isolate or select and study the sensitivity of a specific solution component. First proposed in [25] for linear systems, this framework has since been widely extended to various problems, including linear least squares [26], weighted least squares [27], indefinite least squares [28, 29], total least squares [30], and generalized SPPs [20].
@@
  • In this work, we explore a general form of partial CN, referred to as unified partial CN, which has a versatile nature and provides a comprehensive framework encompassing the partial NCN, MCN, and CCN of the solution of the DSPP.
@@ 2.2 Preliminaries
$\left\|\xi^{\ddagger}\odot(\Upsilon(\bm{v}+{\Delta}\bm{v})-\Upsilon(\bm{v}))\right\|_{\gamma}\leq\mathfrak{K}_{\Upsilon}(\bm{v})\|\chi^{\ddagger}\odot{\Delta}\bm{v}\|_{\tau}\,+\leavevmode\nobreak\ \mathcal{O}(\|\chi^{\ddagger}\odot{\Delta}\bm{v}\|^{2}_{\tau}).$ (2.11)
@@ 3 Partial Unified CNs for the DSPP
In the following theorem, we provide a compact and closed-form expression for the partial unified CN.
@@
Remark 4.
Theorem 2 presents two equivalent expressions (3.6) and (3.7) of the partial NCN. Note that the sizes of the matrices present in these expressions are $k\times(\bm{l}+n^{2}+m^{2}+p^{2}+nm+mp)$ and $k\times k$, respectively. Hence, the expression in (3.7) significantly reduces the storage requirements. Moreover, equation (3.6) contains Kronecker products, which significantly increase memory requirements and computational complexity, making it less practical for large-scale problems. In contrast, equation (3.7) avoids Kronecker products and is more efficient, especially when $k<\bm{l}$, since it results in smaller matrix sizes. Further, to improve computational efficiency and avoid computing $\mathfrak{B}^{-1}$ explicitly present in the expressions of Theorem 2, we adopt a practical approach inspired by the procedure outlined in [34]. For (3.6), we first solve $\mathfrak{B}X=\left[\begin{matrix}\Psi\mathcal{G}&-\bm{\chi}I_{\bm{l}}\end{matrix}\right]$ by using LU decomposition, and then compute $\mathrm{L}X.$ For the expression in (3.7), we first solve the system $\mathfrak{B}Y={\Psi}^{2}\mathcal{J}+{\bm{\chi}}^{2}I_{\bm{l}},$ and following that solve the system $\mathfrak{B}^{{\bm{\top}}}Z=Y,$ and compute $\mathrm{L}Z\mathrm{L}^{{\bm{\top}}}.$ Note that we need to perform only one LU decomposition of $\mathfrak{B}.$
In the next result, to address the computational challenges mentioned in Remark 4 associated with the expressions in Theorem 2, we provide an easily computable upper bound for the partial NCN $\mathfrak{K}^{(2)}_{\Phi}(\mathbf{H},{\bf b};\mathrm{L}).$
@@
In the following result, we provide sharp, computationally less expensive and Kronecker product free upper bounds for the partial MCN and CCN obtained in Theorem 3.
@@ 4 Structured partial CNs
Consider three subspaces $\mathbb{S}_{1}\subseteq{\mathbb{R}}^{n\times n}$, $\mathbb{S}_{2}\subseteq{\mathbb{R}}^{m\times m}$ and $\mathbb{S}_{3}\subseteq{\mathbb{R}}^{p\times p}$ consisting of three distinct linear structured matrices, such as symmetric. Suppose that the corresponding dimensions of the linear subspaces are $s,$ $r$, and $q,$ respectively. Let $A\in\mathbb{S}_{1},$ $D\in\mathbb{S}_{2}$ and $E\in\mathbb{S}_{3},$ then according to [35, 36, 22], there exist unique generating vectors $\bm{a}\in{\mathbb{R}}^{s}$ $\bm{d}\in{\mathbb{R}}^{r}$ and $\bm{e}\in{\mathbb{R}}^{q}$ such that
@@ 6 Numerical experiments
In this part, we present some numerical examples to verify the reliability of the derived partial NCN, MCN, and CCN and their upper bounds for the DSPP. Additionally, we demonstrate their effectiveness in providing tight upper bounds for the relative forward error in solving the DSPP.
@@
respectively.
Similar to the relation in (2.11) and from Definition 2, the forward error can be easily bounded as follows:
$\left\|\xi^{\ddagger}_{\mathrm{L}}\odot\left(\bm{\varphi}(\mathbf{H}+{\Delta}\mathbf{H},\mathbf{b}+{\Delta}\mathbf{b})-\bm{\varphi}(\mathbf{H},{\bf b})\right)\right\|_{\gamma}\leq\bm{\epsilon}\,\mathfrak{K}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})\leavevmode\nobreak\ +\leavevmode\nobreak\ \mathcal{O}(\bm{\epsilon}^{2}).$ (6.1)
Define the following quantities:
@@
Then using the bound in (6.1) and using the setup of NCN, MCN and CCN, $\bm{\epsilon}_{1}\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}),$ $\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ and $\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ can be employed to estimate the upper bound for relative forward errors $\bm{r}_{k},$ $\bm{r}_{m}$ and $\bm{r}_{c},$ respectively. We select the matrix $\mathrm{L}$ as
@@
where $J=\frac{1}{(q+1)^{2}}\,\mathtt{tridiag}(-1,2,-1)\in{\mathbb{R}}^{q\times q},$ $Z=\frac{1}{q+1}\,\mathtt{tridiag}(0,1,-1)\in{\mathbb{R}}^{q\times q}$ and $Y=\mathrm{diag}(1,q+1,\ldots,q^{2}-q+1)\in{\mathbb{R}}^{q\times q}.$ Further, we take $D=I_{m}$ and $E=I_{p}.$ The notation $\mathtt{tridiag}(d_{1},d_{2},d_{3})\in{\mathbb{R}}^{q\times q}$ denotes the tridiagonal matrix with subdiagonal entries $d_{1}$, diagonal entries $d_{2}$ and superdiagonal entries $d_{3}$. Here, the dimension of the coefficient matrix $\mathfrak{B}$ of the DSPP is ${\bm{l}}=4q^{2}.$ We take ${\bf b}=\tt{randn}(n,1)\in{\mathbb{R}}^{\bm{l}}$. Further, we consider $\Psi=\|\mathfrak{B}\|_{F}$ and ${\bm{\chi}}=\|{\bf b}\|_{2}.$ We use Theorems 2 and 3 to compute partial CNs and Corollaries 3.1 and 3.2 for their upper bounds. In Table 1, we listed the computed values of $\bm{\epsilon}_{1}$ and $\bm{\epsilon}_{2}$. The numerical results of the different relative forward errors and their upper bounds for different choices for $\mathrm{L}$ and $q=4:2:16$ are presented in Tables 2-5.
Table 1: Values of $\bm{\epsilon}_{1}$ and $\bm{\epsilon}_{2}$ for different $q.$
$q$ $\bm{\epsilon}_{1}$ $\bm{\epsilon}_{2}$
$4$ $081.1234E-$ 08 $083.2043E-$ 08
$6$ $081.0007E-$ 08 $083.4305E-$ 08
$8$ $099.2718E-$ 09 $082.9944E-$ 08
$10$ $081.0404E-$ 08 $084.0272E-$ 08
$12$ $081.0384E-$ 08 $083.8205E-$ 08
[…]
Table 2: Comparison of relative forward errors and their upper bounds obtained using the NCN, MCN, and CCN, and their upper bounds for $\mathrm{L}=\mathrm{L}_{0}$ for Example 6.1.
$q$ $\bm{r}_{k}$ ${U_{\bm{r}_{k}}}$ ${\widehat{U}_{\bm{r}_{k}}}$ $\bm{r}_{m}$ ${U_{\bm{r}_{m}}}$ ${\widehat{U}_{\bm{r}_{m}}}$ $\bm{r}_{c}$ ${U_{\bm{r}_{c}}}$ ${\widehat{U}_{\bm{r}_{c}}}$
@@
Table 3: Comparison of relative forward errors and their upper bounds obtained using the NCN, MCN, and CCN, and their upper bounds $\mathrm{L}=\mathrm{L}_{n}$ for Example 6.1.
$q$ $\bm{r}_{k}$ ${U_{\bm{r}_{k}}}$ ${\widehat{U}_{\bm{r}_{k}}}$ $\bm{r}_{m}$ ${U_{\bm{r}_{m}}}$ ${\widehat{U}_{\bm{r}_{m}}}$ $\bm{r}_{c}$ ${U_{\bm{r}_{c}}}$ ${\widehat{U}_{\bm{r}_{c}}}$
@@
Table 4: Comparison of relative forward errors and their upper bounds obtained using the NCN, MCN, and CCN, and their upper bounds for $\mathrm{L}=\mathrm{L}_{m}$ for Example 6.1.
$q$ $\bm{r}_{k}$ ${U_{\bm{r}_{k}}}$ ${\widehat{U}_{\bm{r}_{k}}}$ $\bm{r}_{m}$ ${U_{\bm{r}_{m}}}$ ${\widehat{U}_{\bm{r}_{m}}}$ $\bm{r}_{c}$ ${U_{\bm{r}_{c}}}$ ${\widehat{U}_{\bm{r}_{c}}}$
@@
Table 5: Comparison of relative forward errors and their upper bounds obtained using the NCN, MCN, and CCN, and their upper bounds for $\mathrm{L}=\mathrm{L}_{p}$ for Example 6.1.
$q$ $\bm{r}_{k}$ ${U_{\bm{r}_{k}}}$ ${\widehat{U}_{\bm{r}_{k}}}$ $\bm{r}_{m}$ ${U_{\bm{r}_{m}}}$ ${\widehat{U}_{\bm{r}_{m}}}$ $\bm{r}_{c}$ ${U_{\bm{r}_{c}}}$ ${\widehat{U}_{\bm{r}_{c}}}$
@@
${U_{\bm{r}_{k}}}=\bm{\epsilon}_{1}\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}),\leavevmode\nobreak\ {\widehat{U}_{\bm{r}_{k}}}=\bm{\epsilon}_{1}\mathfrak{K}^{(2),u}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}),\leavevmode\nobreak\ {U_{\bm{r}_{m}}}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}),$
@@
${\widehat{U}_{\bm{r}_{m}}}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty,u}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}),\leavevmode\nobreak\ {U_{\bm{r}_{c}}}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}),\leavevmode\nobreak\ {\widehat{U}_{\bm{r}_{c}}}=\bm{\epsilon}_{2}\mathfrak{K}^{\infty,u}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L}).$
By (2.11), the quantities $U_{\bm{r}_{k}}$ and $\widehat{U}_{\bm{r}_{k}}$ are the upper bounds of the relative forward error $\bm{r}_{k}$, $U_{\bm{r}_{m}}$ and $\widehat{U}_{\bm{r}_{m}}$ are the upper bounds of the relative forward error $\bm{r}_{m}$, and $U_{\bm{r}_{c}}$ and $\widehat{U}_{\bm{r}_{c}}$ are the upper bounds of the relative forward error $\bm{r}_{c}.$ These bounds are obtained using the NCN, MCN, CCN, and their upper bounds, respectively.
As by multiplying the same perturbation magnitude $\bm{\epsilon}_{1}$ with the NCN and its upper bound we obtain $U_{\bm{r}_{k}}$ and $\widehat{U}_{\bm{r}_{k}}$, respectively, and as shown in Tables 2–5, they consistently remain of the same order across all values of $q$. Moreover, $U_{\bm{r}_{k}}$ being slightly smaller than $\widehat{U}_{\bm{r}_{k}}$. This demonstrates that the upper bound of the NCN provides a sharp estimate of the NCN. By a similar argument, we found that the upper bounds of the MCN and CCN are also sharp estimates of the MCN and CCN.